# SciClaims: An End-to-End Generative System for Biomedical Claim Analysis

### v1

@@ Abstract
Validating key claims in scientific literature, particularly in biomedical research, is essential for ensuring accuracy and advancing knowledge. This process is critical in sectors like the pharmaceutical industry, where rapid scientific progress requires automation and deep domain expertise. However, current solutions have significant limitations. They lack end-to-end pipelines encompassing all claim extraction, evidence retrieval, and verification steps; rely on complex NLP and information retrieval pipelines prone to multiple failure points; and often fail to provide clear, user-friendly justifications for claim verification outcomes. To address these challenges, we introduce SciClaims, an advanced system powered by state-of-the-art large language models (LLMs) that seamlessly integrates the entire scientific claim analysis process. SciClaims outperforms previous approaches in both claim extraction and verification without requiring additional fine-tuning, setting a new benchmark for automated scientific claim analysis.
@@ 1 Introduction
Systematic Literature Review (SLR) is a crucial field of activity in the biomedical and pharmaceutical industries, as well as in hospital settings. It helps identify clinical data on treatment safety and performance, assess the state of the art, and support regulatory and clinical decision-making. In the pharmaceutical industry, SLR supports activities such as R&D, business development, and medical affairs, while in hospitals, it confirms best practices and guides clinical decisions. By enabling access to relevant information, SLR also shortens regulatory review cycles and facilitates the faster market entry of better treatments.
A key task in SLR is to verify that claims made in scientific texts are supported by previously published, peer-reviewed statements. This process helps define a clear information perimeter for a given work, increasing confidence in the scientific literature. Beyond its role in validating research, the verification of scientific claims also plays a crucial role in key business areas such as detecting patent infringement. In the pharmaceutical industry, for example, companies need to analyze prior scientific publications and patents to defend intellectual property rights. This is particularly relevant in high-stakes legal disputes, such as the recent litigation between Pfizer and BioNTech against Moderna over the mRNA COVID-19 vaccine. [^1] https://iclg.com/news/22358-moderna-wins-german-covid-19-patent-infringement-battle
In this work, we address some of the main challenges that prevent scientific claim analysis systems from becoming a practical solution in such competitive markets as SLR and the pharmaceutical industry, including low accuracies, limited throughput, multi-staged complex pipelines with multiple possible failure points, and lack of user-friendly explanations. In doing so, we look into the different steps involved in the scientific claim analysis pipeline, ranging from the extraction and generation of claims from the original text to the retrieval of potentially relevant evidence to validate those claims and the actual verification against the evidence, supporting or refuting them. Based on state-of-the-art large language models (LLMs), our system SciClaims [^2] SciClaims demo (user/pass: guest/aclindustry2025) available at https://labdemos.expertcustomers.ai/health_claims [^3] Code at: https://github.com/anonymous-sciclaims approaches the entire scientific claim analysis pipeline end-to-end, outperforming previous state-of-the-art methods in both scientific claim extraction and verification without requiring additional fine-tuning.
@@ 2 Related Work
The task of analyzing scientific claims from real-world texts based on background knowledge consists of three primary components: claim extraction, evidence retrieval from a document corpora, and verifying or fact-checking the claims against the evidence Eldifrawi et al. (2024); Vladika and Matthes (2023). However, solving this entire pipeline in an end-to-end manner across the three stages is a challenge that still needs to be met.
@@
Several studies have focused on the challenge of claim extraction from texts. Some frameworks based on zero-shot models Pan et al. (2021); Wright et al. (2022) have achieved promising results, focusing on the generation of datasets of claims from raw texts to train fact-checking models on specific domains. Recent approaches based on LLMs employ these models to extract atomic facts Min et al. (2023); Chern et al. (2023), which serve as summaries of the texts from which they are derived.
@@
For the evidence retrieval phase, dense passage retrieval methods, such as ColBERT Khattab and Zaharia (2020), have emerged due to their ability to retrieve highly relevant documents from large corpora with great precision. However, these methods are computationally expensive and therefore less practical for real-time, lightweight applications. Thus, simpler approaches such as BM-25Robertson and Zaragoza (2009) and Elasticsearch, known for their balance between retrieval quality and computational efficiency, are still preferred for deployment in production environments.
@@
The claim verification component of the pipeline has seen notable advancements as well. The emergence of several claim analysis datasets has resulted in numerous frameworks and models that try to solve this challenge. VERT5ERINI Pradeep et al. (2021), PARAGRAPHJOINT Li et al. (2021), and MultiVerS Wadden et al. (2022) have shown promising results in scientific claim verification datasets such as SciFact Wadden et al. (2020) or CLIMATE-FEVER Diggelmann et al. (2021). Furthermore, due to the recent progress in text generation using large language models (LLMs), a significant body of work has been dedicated to the verification of these artificially generated texts. Frameworks such as FactScore Min et al. (2023), FacTool Chern et al. (2023), or LLM Oasis Scirè et al. (2024) focus on fact-checking automatically generated texts rather than existing human-written passages. However, by leveraging their underlying mechanisms, such as the extraction of atomic knowledge units and claim verification through LLMs, these approaches can also be extended to process scientific documents.
@@
Although individual components have seen substantial progress, an integrated system capable of seamlessly connecting these steps remains an open problem. Many existing systems, such as FactDetect Jafari and Allan (2024) or CliVER Liu et al. (2024) focus on claims that are already identified, neglecting the crucial first step of extracting relevant claims from raw, real-world texts. As a result, these systems are limited to pre-identified claims rather than addressing the full pipeline of claim extraction, retrieval, and verification.
Given the complexity of this problem, our approach seeks to optimize each stage of the pipeline for practical deployment. In particular, we aim to refine the claim extraction, retrieval, and verification modules in a manner that ensures both efficiency and accuracy for real-world applications. By drawing on the strengths of existing systems and focusing on improving their integration, we propose a comprehensive solution for the analysis of biomedical claims in real world texts.
3 Architecture
We present a unified and lightweight architecture built around a Llama3 7B model et al. (2024) and an Elasticsearch index with a verification dataset containing peer-reviewed health and biomedicine literature. This architecture is designed to be lightweight enough to run efficiently on a single 24GB VRAM GPU, enabling high-throughput processing through inference frameworks such as vLLM Kwon et al. (2023), which provide exceptional computation speeds. A seamless processing workflow (see Figure 1) is supported, from claim extraction to evidence retrieval and verification. Rationales in natural language are also provided, which justify system outputs. Next, we present the main building blocks of SciClaims.
@@ 3 Architecture
Claim extraction: This first module extracts potential claims from the source text. A claim is characterized by a specific set of properties detailed in section 5. The Claim Extraction module calls the Llama3 7B model twice: first, to generate an initial list of claims, and second, to refine and filter them, improving the quality of the resulting claims. All the prompts used in our pipeline are provided in Appendix A. Based on our evaluation, we made adjustments to the initial prompts to improve claim extraction performance. Further details on this refinement can be found in section 5.
@@
Document retrieval: The second module retrieves potentially relevant documents from the verification dataset, using the claim as a query to the Elasticsearch index. The verification dataset contains 4.7 million abstracts from PubMed (2000–2022) that were curated using theSemantic Scholar’s Highly Influential Citations metric444www.semanticscholar.org/faq/influential-citations, ensuring that each article is backed by at least three highly influential citations. This selection criterion helps prioritize documents that have been extensively referenced in the academic community, enhancing the quality and relevance of retrieved information for verification. We chose not to filter documents based on Elasticsearch’s scoring mechanism to maximize recall. Instead, the subsequent verification module is responsible for discarding irrelevant documents.
@@
4 System implementation
In this section, we demonstrate the functionality of our approach through a web-based claim analysis tool that allows users to easily interact with the system and analyze claims within relevant texts, leveraging the backend architecture presented in the previous section. The user interface provides a drop-down menu featuring over 30 pre-selected examples from various domains, including biomedicine papers, COVID-related news, social media, and patents (see Appendix B). These examples were chosen to represent a broad spectrum of platforms and relevant disciplines, allowing users to explore diverse contexts. By selecting any example from the list, the corresponding text will be displayed in the text box below, with a hyperlink to the original source. The text box is fully editable, enabling users to modify the content and analyze their own text. While the application can process texts up to 10,000 characters, we recommend keeping the input under 2,000 characters for faster results.
As shown in Appendix B, once the user has selected or entered a text, they can click the Analyze button to initiate the claim analysis process. Results are presented as a list of identified claims, each of which can be expanded to reveal detailed information. For each claim, the interface shows:
  • Prediction label: Whether the claim is supported (green) or refuted (red). Results labeled as Not Enough Information by the LLM are not be presented to the user.
@@ 5 Experimentation and Results
We evaluate our system following a two-staged approach. First, we compare the individual modules on different standard datasets. Next, we take a more holistic approach and evaluate the performance of the system at every step of the pipeline using a large language model (LLM) as a judge.
5.1 Evaluation with datasets
We evaluate the two main modules of the system (claim extraction and claim verification) against curated datasets to assess the performance of the LLMs for each task in isolation.
@@ 5.1.1 Claim extraction
For the evaluation of the claim extraction module, we selected three datasets: SciFact Wadden et al. (2020), FEVER Thorne et al. (2018), and RoSE Liu et al. (2023). SciFact consists of claims derived from scientific articles, paired with evidence from the source document. We use manually crafted claims and their source paragraphs as ground truth to evaluate factual claim extraction. FEVER, which contains general knowledge claims with Wikipedia-based evidence, helps assess performance on non-specialized texts, using Wikipedia passages as ground truth. Since these datasets highlight only selected portions of source content, we compare baseline-generated claims to ground truth using the Levenshtein distance metric, selecting pairs with a similarity score of 0.3 or higher, where such threshold was empirically determined. Additionally, RoSE, originally designed for summarization, helps us to evaluate how well our system synthesizes source abstracts into structured, factually accurate claims.
Method FEVER RoSE SciFact Sentence tokenizer 0.6549 0.6667 0.7211 Pan et al 0.7147 0.7128 0.6561 Wright et al 0.7023 0.6786 0.6475 Noun Phrases Gen 0.7018 0.6889 0.6693 SciClaims 0.7028 0.7184 0.725
Table 1: Similarity Scores in Claim Generation
@@
To assess the performance of the claim generation methods we compute a similarity score using a DeBERTa model fine-tuned on the STS-B dataset [^5] https://huggingface.co/bobox/DeBERTa2-0.9B-ST-stsb. We compare the most similar generated claim with each ground truth claim extracted from the same source paragraph. If there is not any ground truth claim that surpasses our Levenshtein threshold for a given generated claim, the claim is not taken into account. Table 1 shows that SciClaims achieves the highest score on SciFact and RoSE, outperforming the baselines. However, SciClaims is slightly outperformed by Pan et al. (2021) in FEVER, suggesting that our prompt may need additional refinement to generalize to broader domains. Appendix C reports more extensive results.
@@ 5.1.2 Claim verification
The verification task involves predicting the veracity of claim-evidence pairs, classifying each as SUPPORT, REFUTE, or NOT ENOUGH INFORMATION (NEI). Thus, this module is evaluated using the FEVER Thorne et al. (2018) and SciFact Wadden et al. (2020) datasets, where annotated claims-evidence pairs serve as ground truth. Additionally, we use CLIMATE-FEVER Diggelmann et al. (2021), a domain-specific dataset focusing on climate science.
Architecture FEVER CLIMATE-FEVER SciFact Roberta-base* 0.9147 0.4549 0.522 MultiVerS* 0.9381 0.6371 0.7303 SciClaims 0.6467 0.5963 0.6788
Table 2: F1-Scores in Claim Verification. RoBERTa-base and MultiVerS are fine-tuned models, while SciClaims is zero-shot.
To evaluate claim verification we compare our system against two baselines: A fine-tuned RoBERTa-base model Liu et al. (2019) and MultiVerS Wadden et al. (2022), each trained on the respective objective dataset. We evaluate Scifact on its test set, FEVER on its "labelled dev" subset, and we randomly split the CLIMATE-FEVER dataset to follow a standard subset distribution (80% train, 10% validation and 10% test). On these datasets, the fine-tuned models outperform our zero-shot, LLM-based system SciClaims, especially on the FEVER dataset (see Table 2). However, the performance gap significantly narrows on domain-specific datasets like SciFact and CLIMATE-FEVER, highlighting the potential of SciClaims in specialized domains without requiring task-specific training.
@@ 5.2 Evaluation with a judge model
The second stage of our evaluation provides a more comprehensive assessment of the entire system using a large language model as judge. Based on the Judge Arena leaderboard [^6] https://huggingface.co/spaces/AtlaAI/judge-arena, we select Qwen 2.5 72B turbo as judge, since it is the first ranked model with open weights and a higher parameter count than our system’s LLM (Llama3 7B). We chose its 4-bit quantization version (Qwen2.5 72B AWQ777huggingface.co/Qwen/Qwen2.5-72B-Instruct-AWQ) due to hardware limitations. For the evaluation sample, we randomly select 120 documents from the PubMed dataset presented in section 3. The evaluation is conducted in three phases: Claim quality, document retrieval, and claim verification and full-system evaluation.

### v2

@@ Abstract
We present SciClaims, an interactive web-based system for end-to-end scientific claim analysis in the biomedical domain. Designed for high-stakes use cases such as systematic literature reviews and patent validation, SciClaims extracts claims from text, retrieves relevant evidence from PubMed, and verifies their veracity. The system features a user-friendly interface where users can input scientific text and view extracted claims, predictions, supporting or refuting evidence, and justifications in natural language. Unlike prior approaches, SciClaims seamlessly integrates the entire scientific claim analysis process using a single large language model, without requiring additional fine-tuning. SciClaims is optimized to run efficiently on a single GPU and is publicly available for live interaction [^1] https://labdemos.expertcustomers.ai/health_claims (user: guest / password: acldemos2025).
@@ 1 Introduction
Systematic Literature Review (SLR) plays a critical role in biomedical research and the pharmaceutical industry, supporting clinical decisions, regulatory submissions, and R&D pipelines. A central task in SLR is the validation of scientific claims, ensuring that assertions made in scientific texts are supported by prior peer-reviewed research. However, this task is labor-intensive, prone to errors, and increasingly difficult to scale as the number of scientific publications grows.
We present SciClaims, a fully automated system that addresses this challenge by providing an end-to-end scientific claim analysis pipeline in an interactive, user-friendly interface. The system extracts factual claims from scientific texts, retrieves evidence from a curated biomedical corpus, and verifies the validity of each claim using large language models (LLMs). To improve transparency, SciClaims also offers natural language rationales and highlights key supporting or refuting evidence.
SciClaims is optimized for real-world deployment and operates efficiently on a single GPU, supports high-throughput processing, and handles documents of up to 10,000 characters. It is accessible through a web-based interface, allowing users to analyze both preloaded and custom input texts.
In this demonstration, we showcase SciClaims as a useful tool to enable users to validate scientific claims in real time, facilitating trustworthy knowledge discovery in high-stakes domains like biomedicine and pharmaceuticals. A how-to video222www.youtube.com/watch?v=jyms_Ey0YSQ and all the code333www.github.com/expertailab/sciclaims-backend and www.github.com/expertailab/sciclaims-frontend used in this project have been published and made publicly available.
@@ 2 Related Work
The task of analyzing scientific claims from real-world texts based on background knowledge consists of three primary components: claim extraction, evidence retrieval from a document corpus, and verifying or fact-checking the claims against the evidence Eldifrawi et al. (2024); Vladika and Matthes (2023). However, solving this pipeline end-to-end across all three stages remains an open challenge.
@@
Several studies have addressed the challenge of extracting claims from scientific texts. Some frameworks based on zero-shot models Pan et al. (2021); Wright et al. (2022) have achieved promising results, primarily focusing on generating claim datasets from raw texts to train fact-checking models in specific domains. However, these methods rely on multi-stage NLP pipelines that are prone to failure and tend to produce a large number of claims per document, making their integration into an end-to-end system difficult. In contrast, recent approaches based on LLMs extract atomic factual units Min et al. (2023); Chern et al. (2023), which serve as concise, interpretable summaries of the source texts.
@@
For the evidence retrieval phase, dense passage retrieval methods, such as ColBERT Khattab and Zaharia (2020), have emerged due to their ability to retrieve highly relevant documents from large corpora with great precision. However, these methods are computationally expensive and therefore less practical for real-time, lightweight applications. Thus, simpler approaches such as BM-25 Robertson and Zaragoza (2009) and tools like Elasticsearch, known for their balance between retrieval quality and computational efficiency, are still preferred for deployment in production environments.
@@
The release of several claim verification datasets has spurred the development of numerous models aimed at addressing this challenge. VERT5ERINI Pradeep et al. (2021), PARAGRAPHJOINT Li et al. (2021), and MultiVerS Wadden et al. (2022) have achieved promising results on scientific benchmarks like SciFact Wadden et al. (2020) and CLIMATE-FEVER Diggelmann et al. (2021). However, these approaches still leave room for improvement and typically rely on large, labeled datasets, which limits their scalability across domains. In parallel, recent advances in LLMs have led to increased interest in verifying automatically generated content. Frameworks such as FactScore Min et al. (2023), FacTool Chern et al. (2023), and LLM Oasis scirè2024truthmirageendtoendfactuality focus on assessing the factuality of LLM-generated text rather than existing, human-written passages. Nonetheless, their techniques, such as extracting atomic knowledge units and using zero-shot LLMs for claim extraction and verification, can be extended to real-world scientific texts, offering a promising direction for scalable, data-efficient claim analysis.
@@
Although individual components have seen substantial progress, an integrated system capable of seamlessly connecting these steps remains an open problem. Many existing systems, such as FactDetect Jafari and Allan (2024), CliVER Liu et al. (2024), and more recently Liang and Sonntag (2025); Vladika et al. (2025), focus on claims that are already identified, neglecting the crucial first step of extracting relevant claims from raw, real-world texts. As a result, these systems are limited to pre-identified claims rather than addressing the full pipeline of claim extraction, retrieval, and verification.
Given the complexity of this problem, our approach aims to optimize each stage of the pipeline to ensure both efficiency and accuracy in real-world biomedical claim analysis. In this work, we tap on the strengths of modern LLMs to address previous limitations, specifically in the claim extraction and verification stages, and integrate our results as a robust online system.
3 System Description
Figure 2: Screenshot of the SciClaims Demo.
SciClaims is an interactive, end-to-end system for scientific claim analysis, comprising three main components: claim extraction, evidence retrieval, and claim verification. It is optimized to run efficiently on a single 24GB VRAM GPU, enabling real-time performance via a web-based interface.
The system architecture (Figure 1) is built around a Llama3 8B Instruct model Grattafiori et al. (2024) and an Elasticsearch-based retrieval engine. It processes biomedical or scientific text input and outputs a list of extracted claims, their verification status, relevant evidence, and a natural language rationale. The model is set up using vLLM Kwon et al. (2023), which enables high-throughput processing in inference. Next, we present the main building blocks of SciClaims.
@@ 3 System Description
Claim extraction: This first module extracts potential claims from the source text. A claim is characterized by a specific set of properties detailed in section 5. The Claim Extraction module calls the Llama3 8B Instruct model twice: first, to generate an initial list of claims, and second, to refine and filter them, improving the quality of the resulting claims. All the prompts used in our pipeline are provided in Appendix A. Based on our evaluation, we made adjustments to the initial prompts to improve claim extraction performance. Further details on this refinement can be found in section 5.
@@
Document retrieval: The second module retrieves potentially relevant documents from the verification dataset, using the claim as a query to the Elasticsearch index. The verification dataset contains 4.7 million abstracts from PubMed (2000–2022) that were curated using the Semantic Scholar’s Highly Influential Citations metric 1, ensuring that each article is backed by at least three highly influential citations. This selection criterion helps prioritize documents that have been extensively referenced in the academic community, enhancing the quality and relevance of retrieved information for verification. We chose not to filter documents based on Elasticsearch’s scoring mechanism to maximize recall. Instead, the subsequent verification module is responsible for discarding irrelevant documents.
@@
4 User interface
In this section, we demonstrate the functionality of our approach through a web-based claim analysis tool that allows users to easily interact with the system and analyze claims within relevant texts, leveraging the backend architecture presented in the previous section. The demo showcases the complete workflow, from entering a scientific passage to obtaining fact-checking results with supporting evidence and explanations.
The user interface provides a drop-down menu featuring over 30 pre-selected examples from various domains, including biomedicine papers, COVID-related news, social media, and patents (see Figure 2). These examples were chosen to represent a broad spectrum of platforms and relevant disciplines, allowing users to explore diverse contexts. By selecting any example from the list, the corresponding text will be displayed in the text box below, with a hyperlink to the original source. The text box is fully editable, enabling users to modify the content and analyze their own text. While the application can process texts up to 10,000 characters, we recommend keeping the input under 2,000 characters for faster results.
As shown in Figure 2, once the user has selected or entered a text, they can click the Analyze button to initiate the claim analysis process. Results are presented as a list of identified claims, each of which can be expanded to reveal detailed information. For each claim, the interface shows:
Prediction label: Whether the claim is supported (green) or refuted (red). Results labeled as Not Enough Information by the LLM are not presented to the user.
@@ 4 User interface
This interactive interface provides an intuitive and user-friendly environment for testing and exploring claim analysis in various health-related texts. The analysis goes through the entire pipeline, offering a label for the claims extracted along with the retrieved evidence. When a claim receives conflicting labels from different pieces of evidence, our system returns all relevant pairs, allowing users to assess which is more accurate. Since scientific knowledge evolves, evidence considered true at one time may later be contradicted. We recognize the value of providing all supporting and refuting evidence along with their publication dates.
@@ 5 Experimentation and Results
We evaluate our system following a three-stage approach. First, we evaluate SciClaims’ modules against some baselines using SciFact as benchmark. Then, we evaluate the performance of the system at every step of the pipeline using an LLM as a judge. Finally, we perform a human evaluation for the whole system.
5.1 Evaluation in SciFact
We use SciFact, a benchmark dataset focused on biomedical literature, to evaluate SciClaims. It includes 1,400 expert-authored scientific claims, each linked to evidence-containing abstracts and annotated with a verification label.
SciFact supports two complementary evaluations: (1) claim extraction, by comparing system-generated claims to human-written ones using shared source documents, and (2) claim verification, by assessing how accurately the system labels claim-evidence pairs.
@@ 5.1.1 Claim extraction
To evaluate the claim extraction module, we consider the expert-written claims from SciFact as the ground truth for each source paragraph.
Method Rouge1 Rouge2 RougeL Similarity Score Sentence tokenizer 0.3313 0.1980 0.3030 0.7211 Pan et al. (2021) 0.2780 0.1334 0.2555 0.6561 Wright et al. (2022) 0.2525 0.0762 0.2220 0.6475 Noun Phrases Gen 0.2567 0.0953 0.2293 0.6693 SciClaims 0.3387 0.1896 0.3084 0.7250
Table 1: Claim quality scores in SciFact
@@
We evaluate the quality of claim generation by matching each system-generated claim with its most similar gold claim from the same paragraph. We discard pairs with a Levenshtein similarity score below 0.3, a threshold empirically tuned to balance recall and noise. For valid matches, we compute ROUGE-1, ROUGE-2, ROUGE-L, and a semantic similarity score using a DeBERTa model fine-tuned on the STS-B dataset. As shown in Table 1, SciClaims achieves the highest scores in ROUGE-1, ROUGEL, and semantic similarity, indicating more faithful and informative claim generation.
@@ 5.1.2 Claim verification
We evaluate SciClaims on SciFact’s test set, comparing its label accuracy for claim-evidence pairs with existing approaches.
Architecture Precision Recall F1-Score Roberta-base* 0.4662 0.5963 0.5220 MultiVerS* 0.7286 0.7321 0.7303 SciClaims (LLaMA3 8B Instruct) 0.7034 0.6863 0.6788 SciClaims (Qwen2 7B Instruct) 0.6649 0.6677 0.6439 SciClaims (Phi3 Small 8K) 0.7100 0.6894 0.6525 SciClaims (OLMO2 7B Instruct) 0.6379 0.6118 0.5886
Table 2: Precision, recall and F1-Scores in Claim Verification of SciFact test set. RoBERTa-base and MultiVerS are fine-tuned models, while SciClaims is zero-shot. In parenthesis, the backbone used in each SciClaims configuration
We compare SciClaims with two strong baselines: a RoBERTa-base classifier Liu et al. (2019) and MultiVerS Wadden et al. (2022), a longformer-based model for claim verification. Both baselines are fine-tuned on SciFact. While MultiVerS achieves the highest F1-score, SciClaims performs competitively despite operating in a zero-shot setting. As shown in Table 2, the relatively narrow performance gap demonstrates SciClaims’ strong generalization ability without task specific training. Furthermore, as shown in Table 2, we evaluate SciClaims with different LLMs with similar sizes as backbone, being LLaMA3 8B Instruct the one which offered the best performance.
@@ 5.2 Evaluation with a judge model
The second stage of our evaluation provides a comprehensive assessment of the entire system using a LLM as judge. Based on the Judge Arena leaderboard [^4] https://huggingface.co/spaces/AtlaAI/judge-arena, we select Qwen 2.5 72B turbo as judge, since it is the first ranked model with open weights and a higher parameter count than our system’s LLM (Llama3 8B Instruct). We chose its 4-bit quantization version (Qwen2.5 72B AWQ555huggingface.co/Qwen/Qwen2.5-72B-Instruct-AWQ) due to hardware limitations. For the evaluation sample, we randomly select 120 documents from the PubMed dataset presented in section 3. The evaluation is conducted in three phases: Claim quality, document retrieval, and claim verification and full-system evaluation.